# Are Generative Models Underconfident? Better Quality Estimation with Boosted Model Probability

### v2

@@ Abstract
Quality Estimation (QE) is estimating quality of the model output during inference when the ground truth is not available. Deriving output quality from the models’ output probability is the most trivial and low-effort way. However, we show that the output probability of text-generation models can appear underconfident. At each output step, there can be multiple correct options, making the probability distribution spread out more. Thus, lower probability does not necessarily mean lower output quality. Due to this observation, we propose a QE approach called BoostedProb, which boosts the model’s confidence in cases where there are multiple viable output options. With no increase in complexity, BoostedProb is notably better than raw model probability in different settings, achieving on average +0.194 improvement in Pearson correlation to ground-truth quality. It also comes close to or outperforms more costly approaches like supervised or ensemble-based QE in certain settings.
@@ 1 Introduction
Particularly, Quality Estimation (QE) is the task of providing quality scores of model outputs during inference when the ground truth is not available. The most straightforward way is to utilize the model’s output probability. While previous works have shown that model probability is prone to be overconfident Nguyen et al. (2015); Li et al. (2021), in this work, we point out another issue. We show that the output probability on free-form text generation tasks, such as translation or summarization, can be underconfident. Specifically, lower probability does not necessarily indicate lower output quality, but could mean that the probability distribution is spread out on multiple correct options.
@@
    We propose a QE approach, BoostedProb, that boosts the confidence of these dominant tokens. BoostedProb is easy to implement and does not add any complexity compared to raw model probabilities. It is substantially more efficient than ensemble-based QE which requires generating multiple outputs, and supervised QE which is data-dependent and not available for tasks other than translation.
@@ 2.1 Quality Estimation
One outstanding case of external QE module is supervised QE models for Machine Translation (MT), such as CometKiwi Rei et al. (2022). For MT, there exists abundant data of (source, model translation, human-labeled scores) tuples, which enables training supervised QE models. One can try to avoid the use of costly human-labeled scores by training QE models on synthetic data with synthetic errors Tuan et al. (2021), or synthetic scores using reference-based metrics Zouhar et al. (2023) like BLEU Papineni et al. (2002), chrF Popović (2015) or BERTScore Zhang et al. (2019). Supervised QE has been widely adopted in MT, and is getting close to the performance of reference-based metrics Freitag et al. (2022).
@@ 2.2 Dominant Tokens
Previous works have considered that there can be multiple tokens with dominant probability mass in the output distribution. For example, Ott et al. (2018) shows that, for MT, model distribution is highly spread in the hypothesis space. However, they focus on its effect on model fitting and inference search rather than for QE. Other works focus on sampling, where they try to find the set of dominant tokens to sample from during output generation to maintain high quality but also high diversity. Popular sampling strategies includes top-$k$ Fan et al. (2018), top-$p$ Holtzman et al. , $\epsilon$-cut Hewitt et al. (2022), $\eta$-cut Hewitt et al. (2022) and min-$p$ Nguyen et al. (2024). For top-$k$, the assumption is that, the top $k$ tokens with the highest probability are the most important ones. For top-$p$, the most important tokens are those with top probabilities that sum up to $p$. For $\epsilon$-cut, the most important token probabilities are larger than $\epsilon$. For $\eta$-cut, the most important token probabilities are larger than either $\eta$ or $\sqrt{\eta}\times\exp(-entropy(\mathbb{P}))$, where $\mathbb{P}$ is the output probability distribution. For min-$p$, the most important tokens have probabilities larger than the top-1 probability multiplied by $p$.
In our work, we focus on finding dominant tokens to boost their confidence for QE, rather than to support sampling or inference search.
@@ 3 Model Probability – An Underconfident Quality Estimator?
Our analysis shows that, text-generation tasks with ambiguity introduce aleatoric uncertainty, i.e., uncertainty coming from the data, which differs from epistemic uncertainty, i.e., uncertainty coming from the model’s incompetence. Aleatoric uncertainty makes the model appear underconfident, as the probability mass is spread over multiple valid options. This brings us to a simple modification to the model probability to improve its effectiveness as a quality estimator, as detailed below.
@@ 4 BoostedProb
We arrived at a condition that considers both the relative value (drops larger than $x$%) and absolute value (drops larger than $\epsilon$), making our approach more flexible in finding the dominant tokens in different probability distributions.
@@
where tokens with probabilities above the cutting point are dominant, and others are non-dominant.
Figure 3: A dominant cluster found by our heuristic.
@@
$QE(w_{(i)})=\begin{cases}p_{(i)},&\text{if }i>c\\ \sum\limits_{j=1}^{c}p_{(j)},&\text{otherwise }i\leq c\\ \end{cases}$
@@ 5.3 Baselines
Average 0.092 0.221 0.286
@@ 5.4 Hyperparameters
We loosely tune the hyperparameters, i.e., $x$ and $\epsilon$, on three models: Whisper, DeltaLM Large and Tower (see Appendix B). For these models, $x=30\%$ and $\epsilon=0.005$ are either the best or close to the best set of hyperparameters, showing that our approach is robust to hyperparameter setup. We then use these hyperparameters for all experiments.
@@ 5.5.1 Automatically Generated Gold Quality
We use XCOMET-XL Guerreiro et al. (2024) to generate gold quality for translations. XCOMET-XL is a reference-based neural model. Dinh et al. (2024) showed that, for MT, such reference-based neural metrics are good enough to be used as the ground truth to rank reference-free QE metrics.
@@
We use BART Score Yuan et al. (2021) as gold output quality. The quality scores are calculated as the BART probability of the output given the input text. Unlike for MT, there has not been any study showing that reference-based metrics like Bart Score are sufficient as the ground truth for reference-free QE metrics. Therefore, we additionally report on other reference-based metrics in Appendix C.
@@ 6.1 Overall Performance
We also compare BoostedProb with a more advanced method, Monte Carlo sequence entropy Malinin and Gales (2021); Kuhn et al. (2023), in Table 4. The results are indecisive, as our approach only outperforms this baseline in some settings. However, recall that this baseline is more costly, as we would have to generate multiple output samples.
Model Lang. Monte Boosted
Carlo Prob
MT NLLB en-de 0.303 0.525
zh-en 0.337 0.289
Tower en-de 0.302 0.414
zh-en 0.240 0.240
Sum. Bloomz 560M en 0.236 0.210
Llama3.2 3B en 0.005 0.209
QA Bloomz 560M en 0.003 0.009
Llama3.2 3B en 0.175 0.111
Table 4: BoostedProb vs. Monte Carlo entropy.
@@ 6.2 Scoring Other Models’ Output (Prism)
{adjustwidth}
@@
Table 5: Performance of QE methods with Prism, in Pearson correlation to human-labeled quality score
@@ 6.4 Effect of Generative Performance
We see that the QE performance of the raw model probability stays consistently low. However, with BoostedProb, the QE performance improves as the quality of the model improves. This indicates that, with BoostedProb, improving a model’s quality would come with improving its ability to do self-evaluation. Figure 4 also exposes a limitation of our approach, since it worsens the QE performance compared to raw probability for very weak models. The reason might be due to our approach mistakenly emphasizing the weak models’ overconfidence (Appendix D). In these cases, we recommend using a stronger model to score the output of the weak model, rather than using the probability of the weak model to score itself, like in the Prism setting (Section 6.2).
@@ 6.5 Finding Dominant Cluster
However, as discussed in Section 4, using our method would still be a safer choice, as it would avoid mistakenly considering very low probability tokens as dominant, in the rare case that they are selected as the final output. Additionally, our method is more robust to hyperparameters: in three out of four settings in Table 7, we arrive at $x=30\%,\epsilon=0.005$, which are the same hyperparameter values we found before (Appendix B).
@@ 8 Limitations
BoostedProb does not work well with very weak models, as it might emphasize weak models’ overconfidence. This is mentioned in Section 6.4, and further discussed in Appendix D. In these cases, we recommend using BoostedProb with a strong model to score the output of the weak model. This is the setting of Prism, which we have reported on in Section 6.2. BoostedProb is also unlikely to have an effect for less ambiguous text generation tasks like Automatic Speech Recognition, or multiple-choice Question-Answering, since the dominant clusters with sizes larger than one are unlikely to appear (see Appendix E). One can also argue that, unsupervised QE methods are in general not very useful whenever a supervised model exists (like in Speech Translation and Machine Translation). However, in these cases, our approach still brings benefits in terms of simplicity in implementation and inference time, compared to having an extra module for supervised QE in a translation pipeline.
@@ Appendix A Human Analysis of Valid Output Tokens
Appendix B Hyperparameters Tuning
B.1 Finding Dominant Tokens
We tune the hyperparameters for our approach, i.e., the value of $x$that defines the relative threshold, and the value of $\epsilon$ that defines the absolute threshold that makes a reduction in probability mass significant, thus separating dominant from non-dominant tokens. The set of candidate values for $x$is $\{0.2,0.3,0.4,0.5,0.6\}$. The set of candidate values for $\epsilon$ is $\{0.005,0.01,0.1\}$.
@@ B.1 Finding Dominant Tokens
B.2 Sequence Aggregation of Token Scores
@@
Appendix C Correlation with ROUGE-L and BertScore
@@
Table 9: Correlation of different quality scores to different ground truth metrics: Bart Score, RougeL and BertScore.
@@
Appendix D Negative Effect on Very Weak Models
As mentioned in Section 6.4, BoostedProb improves QE performance of output probability for stronger models, but worsens it for weak models. This is somewhat expected, since the motivation of BoostedProb is to improve cases when the model is underconfident. It does not consider the cases when a low-quality model is overconfident and constantly assigns high probability values to the wrong token. To test whether this is truly the cause, we manually look at some output by the worse-performing model, Whisper Tiny, on Chinese-to-English test data. One example is as follows:
Source: {CJK*}UTF8gbsn "有了它,我们才有了火车、汽车和许多其他交通工具"

### v4

@@ Abstract
Quality Estimation (QE) is estimating the quality of the model output during inference when the ground truth is not available. Deriving output quality from the models’ output probability is the most trivial and low-effort way. However, we show that the output probability of text-generation models can appear underconfident. At each output step, there can be multiple correct options, making the probability distribution spread out more. Thus, lower probability does not necessarily mean lower output quality. Due to this observation, we propose a QE approach called BoostedProb [^1] Implementation available at https://github.com/TuAnh23/boostedprob., which boosts the model’s confidence in cases where there are multiple viable output options. With no increase in complexity, BoostedProb is notably better than raw model probability in different settings, achieving on average +0.194 improvement in Pearson correlation to ground-truth quality. It also comes close to or outperforms more costly approaches like supervised or ensemble-based QE in certain settings.
@@ 1 Introduction
Particularly, Quality Estimation (QE) is the task of providing quality scores of model outputs during inference when the ground truth is not available. The most straightforward way is to utilize the model’s output probability. While previous works have shown that model probability is prone to be overconfident Nguyen et al. (2015); Li et al. (2021), in this work, we point out another issue. We show that the output probability on free-form text generation tasks, such as translation or summarization, can be underconfident. Specifically, lower probability does not necessarily indicate lower output quality, but could mean that the probability distribution is spread out over multiple correct options.
@@
    We propose a QE approach, BoostedProb, that boosts the confidence of these dominant tokens. BoostedProb is easy to implement and does not add any complexity compared to raw model probabilities. It is substantially more efficient than ensemble-based QE, which requires generating multiple outputs, and supervised QE, which is data-dependent and not available for tasks other than translation.
@@ 2.1 Quality Estimation
One outstanding case of external QE modules is supervised QE models for Machine Translation (MT), such as CometKiwi Rei et al. (2022). For MT, there exists abundant data of (source, model translation, human-labeled scores) tuples, which enables training supervised QE models. One can try to avoid the use of costly human-labeled scores by training QE models on synthetic data with synthetic errors Tuan et al. (2021), or synthetic scores using reference-based metrics Zouhar et al. (2023) like BLEU Papineni et al. (2002), chrF Popović (2015) or BERTScore Zhang et al. (2019). Supervised QE has been widely adopted in MT, and is getting close to the performance of reference-based metrics Freitag et al. (2022).
@@ 2.2 Dominant Tokens
Previous works have considered that there can be multiple tokens with dominant probability mass in the output distribution. For example, Ott et al. (2018) shows that, for MT, model distribution is highly spread in the hypothesis space. However, they focus on its effect on model fitting and inference search rather than on QE. Other works focus on sampling, where they try to find the set of dominant tokens to sample from during output generation to maintain high quality but also high diversity. Popular sampling strategies includes top-$k$ Fan et al. (2018), top-$p$ Holtzman et al. , $\epsilon$-cut Hewitt et al. (2022), $\eta$-cut Hewitt et al. (2022) and min-$p$ Nguyen et al. (2024). For top-$k$, the assumption is that, the top $k$ tokens with the highest probability are the most important ones. For top-$p$, the most important tokens are those with top probabilities that sum up to $p$. For $\epsilon$-cut, the most important token probabilities are larger than $\epsilon$. For $\eta$-cut, the most important token probabilities are larger than either $\eta$ or $\sqrt{\eta}\times\exp(-entropy(\mathbb{P}))$, where $\mathbb{P}$ is the output probability distribution. For min-$p$, the most important tokens have probabilities larger than the top-1 probability multiplied by $p$.
In our work, we focus on finding dominant tokens to boost their confidence for QE, rather than to support sampling or search strategies during inference.
@@ 3 Model Probability – An Underconfident Quality Estimator?
Our analysis shows that, text-generation tasks with ambiguity introduce aleatoric uncertainty, i.e., uncertainty coming from the data, which differs from epistemic uncertainty, i.e., uncertainty coming from the model’s incompetence. Aleatoric uncertainty makes the model appear underconfident, as the probability mass is spread over multiple valid options. We discuss this underconfidence phenomenon more formally with a theoretical analysis of the softmax function in Appendix B, where we show that there exists an upperbound of the probability scores assigned to every correct token at an output step, which is dependent on the number of correct tokens. This observation brings us to a simple modification to the model probability to improve its effectiveness as a quality estimator, as detailed below.
@@ 4 BoostedProb
We arrived at a condition that considers both the relative value (drops larger than xx%) and absolute value (drops larger than $\epsilon$), making our approach more flexible in finding the dominant tokens in different probability distributions.
@@
where tokens with probabilities above the cutting point are dominant, and others are non-dominant. An illustration is shown in Figure 3.
@@
[…]
@@
We theoretically show that BoostedProb helps tackle the underconfidence phenomenon discussed in Section 3 by allowing multiple correct output tokens to be assigned with a high, arbitrarily close to 1 score, which we detail in Appendix B.
@@ 5.2 Models
For smaller models, i.e., Whisper and DeltaLM, we generate output using beam search with beam size 4. For other models, we generate output with greedy search.
@@ 5.3 Baselines
Question Answering Bloomz 560M SciEx en,de -0.002 0.005 0.006
(University Exam) Llama3.2 3B SciEx en,de 0.002 0.228 0.310
Llama3.3 70B SciEx en,de 0.103 0.180 0.180
Average 0.083 0.208 0.268
@@
LLM self-judge
We also compare our approach against a recent baseline, LLM-as-a-Judge Zheng et al. (2023). To make this baseline more comparable to our reference-free QE, no-external-model setting, we adapt it to an LLM self-judge setup, where the model is asked to assign a quality score to its own output. This method requires an additional inference step and is not applicable to task-specific models such as Whisper and NLLB.
@@ 5.4 Hyperparameters
We loosely tune the hyperparameters, i.e., $x$ and $\epsilon$, on three models: Whisper, DeltaLM Large and Tower (see Appendix C). For these models, $x=30\%$ and $\epsilon=0.005$ are either the best or close to the best set of hyperparameters, showing that our approach is robust to hyperparameter setup. We then use these hyperparameters for all experiments.
@@ 5.5.1 Automatically Generated Gold Quality
We create pseudo ground-truth quality scores to evaluate our reference-free QE methods using reference-based metrics. Reference-based metrics use human ground-truth answers in order to assign a quality score to a model output. We expect reference-based metrics to produce more reliable quality scores compared to reference-free QE methods, thus choosing them as pseudo ground-truth for evaluation.
@@
We use XCOMET-XL Guerreiro et al. (2024) to generate pseudo ground-truth quality for translations. XCOMET-XL is a reference-based neural model. Dinh et al. (2024b) showed that, for MT, such reference-based neural metrics are good enough to be used as the ground truth to rank reference-free QE metrics.
@@
We use BART Score Yuan et al. (2021) as pseudo ground-truth output quality. The quality scores are calculated as the BART Lewis et al. (2020) model probability of the output given the input text. Unlike for MT, there has not been any study showing that reference-based metrics like Bart Score are sufficient as the ground truth for reference-free QE metrics. Therefore, we additionally report on other reference-based metrics in Appendix E, including RougeL Lin (2004), BertScore Zhang et al. (2019), and LLM-as-a-Judge Zheng et al. (2023) with Qwen2.5 72B Instruct Team (2024).
@@ 6.1 Overall Performance
We also compare BoostedProb with more advanced methods, namely Monte Carlo sequence entropy Malinin and Gales (2021); Kuhn et al. (2023) and LLM self-judge. As shown in Table 4, the results are mixed: our approach outperforms these baselines in certain settings. It is important to note, however, that these baselines are more costly and less flexible. Monte Carlo sequence entropy requires generating multiple output samples, while LLM self-judge requires an additional inference pass and is not applicable to task-specific models.
Model Lang. Monte LLM Self Boosted
Carlo judge Prob
MT NLLB en-de 0.303 - 0.525
zh-en 0.337 - 0.289
Tower en-de 0.302 - 0.414
zh-en 0.240 - 0.240
Sum. Bloomz 560M en 0.236 0.149 0.210
Llama3.2 3B en 0.005 0.159 0.209
QA Bloomz 560M en 0.003 0.019 0.009
Llama3.2 3B en 0.175 0.232 0.111
Table 4: BoostedProb vs. Monte Carlo entropy and LLM self-judge. LLM self-judge is not applicable for task-specific models.1
11footnotetext: Tower is an LLM; however, we often observe that it fails in our specific self-judge setting, i.e., fails to produce a single quality score for a given translation. This is likely because the version of Tower we used was not trained for this task.
One potential concern is whether the differences in the inference process affect the reported results. BoostedProb is designed to be applicable across inference methods (e.g., greedy decoding, beam search, top-p sampling), since it does not rely on the model always selecting the top-probability token. Furthermore, to address potential variability due to randomness, we provide example results across multiple runs with different random seeds in Appendix D.
@@ 6.2 Scoring Other Models’ Output (Prism)
Scoring Model Prism Original NLLB
Scored Model Best Worst Best Worst
MT MT MT MT Avg.
en-de
Probability 0.020 0.130 0.068 0.061 0.070
Entropy 0.056 0.147 0.123 0.318 0.161
@@
Supervised QE 0.202 0.453 0.202 0.453 0.328
zh-en
Probability 0.205 0.285 0.153 0.085 0.182
Entropy 0.246 0.299 0.095 0.194 0.209
[…]
Supervised QE 0.341 0.429 0.341 0.429 0.385
Table 5: Performance of QE methods with Prism, in Pearson correlation to human-labeled quality score.
@@ 6.4 Effect of Generative Performance
We see that the QE performance of the raw model probability stays consistently low. However, with BoostedProb, the QE performance improves as the quality of the model improves. This indicates that, with BoostedProb, improving a model’s quality would come with improving its ability to do self-evaluation. Figure 4 also exposes a limitation of our approach, since it worsens the QE performance compared to raw probability for very weak models. The reason might be due to our approach mistakenly emphasizing the weak models’ overconfidence (Appendix F). In these cases, we recommend using a stronger model to score the output of the weak model, rather than using the probability of the weak model to score itself, like in the Prism setting (Section 6.2).
@@ 6.5 Finding Dominant Cluster
However, as discussed in Section 4, our method would still be a safer choice, as it would avoid mistakenly considering very low probability tokens as dominant, in the rare case that they are selected as the final output. Additionally, our method is more robust to hyperparameters: in three out of four settings in Table 7, we arrive at $x=30\%,\epsilon=0.005$, which are the same hyperparameter values we found before (Appendix C).
@@ 8 Limitations
BoostedProb does not work well with very weak models, as it might emphasize weak models’ overconfidence. This is mentioned in Section 6.4, and further discussed in Appendix F. In these cases, we recommend using BoostedProb with a strong model to score the output of the weak model. This is the setting of Prism, which we have reported on in Section 6.2. BoostedProb is also unlikely to have an effect for less ambiguous text generation tasks like Automatic Speech Recognition, or multiple-choice Question-Answering, since the dominant clusters with sizes larger than one are unlikely to appear (see Appendix G). One can also argue that, unsupervised QE methods are in general not very useful whenever a supervised model exists (like in Speech Translation and Machine Translation). However, in these cases, our approach still brings benefits in terms of simplicity in implementation and inference time, compared to having an extra module for supervised QE in a translation pipeline.
Acknowledgements
This work was supported by the Helmholtz Programme-oriented Funding, with project number 46.24.01, project name AI for Language Technologies. We acknowledge the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Wurttemberg and by the Federal Ministry of Education and Research. This work also received support from the European Union’s Horizon research and innovation programme under grant agreement No 101135798, project Meetween (My Personal AI Mediator for Virtual MEETtings BetWEEN People).
@@ Appendix A Human Analysis of Valid Output Tokens
Appendix B Theoretical Analysis
Language models generate text by predicting the next token $y_{t}$ given the previous context $y_{<t}$ and the input $X$, using conditional probability:
$P(y_{t}|y_{<t},X)=softmax(z_{t})$
where $z_{t}\in R^{|V|}$ is the logit vector at output step $t$ over the vocabulary $V$.
The softmax function defines the probability of the selected token $w_{i}\in V$ as:
$P(y_{t}=w_{i}|y_{<t},X)=\frac{e^{z_{t,i}}}{\sum_{j=1}^{|V|}e^{z_{t,j}}}$
The softmax function satisfies: $P(y_{t}=w_{i})\geq 0$ and $\sum_{i=1}^{|V|}P(y_{t}=w_{i})=1$.
In natural language, multiple tokens can be the valid next output. Let:
  • $C=\{w_{c1},w_{c2},...\}\subset V$: the set of correct tokens at step $t$
  • $|C|=k$: number of correct tokens.
If we want each correct token to have a minimum probability $p_{min}$, then:
$\sum_{i\in C}P(y_{t}=w_{i})\geq k\cdot p_{min}$
However, since $\sum_{i\in V}P(y_{t}=w_{i})=1$, we must have:
$k\cdot p_{min}\leq 1\quad\Rightarrow\quad p_{min}\leq\frac{1}{k}$
Thus, as $k$ increases, the maximum assignable probability to each correct token decreases, leading to underconfidence.
With BoostedProb, we instead use the sum of the probability mass of $C$ as the quality score:
$BooostedProb(y_{t}=w_{c1})$
$=BooostedProb(y_{t}=w_{c2})$
$=P_{C}=\sum_{i\in C}P(y_{t}=w_{i}|y_{<t},X)$
The only condition on $P_{C}$ is that $P_{C}\leq 1$. Therefore, with BoostedProb, $p_{min}$ can have a high value close to one, regardless of the size of $C$, thus tackling the underconfidence issue.
Appendix C Hyperparameters Tuning
C.1 Finding Dominant Tokens
We tune the hyperparameters for our approach, i.e., the value of xx% that defines the relative threshold, and the value of $\epsilon$ that defines the absolute threshold that makes a reduction in probability mass significant, thus separating dominant from non-dominant tokens. The set of candidate values for xx% is $\{0.2,0.3,0.4,0.5,0.6\}$. The set of candidate values for $\epsilon$ is $\{0.005,0.01,0.1\}$.
@@ C.1 Finding Dominant Tokens
C.2 Sequence Aggregation of Token Scores
@@
Appendix D Randomness in Inference
To account for randomness in the inference process, it is ideal to repeat each experiment with different random seeds and report the variance. However, this approach is computationally expensive. In our work, we instead prioritize running a broad set of experiments across models and tasks, rather than repeating each experiment multiple times. For illustration, we conduct an additional study with multiple seeds in one setting: the NLLB model on a translation task. Specifically, we apply top-$p$ sampling with $p=0.5$ and repeat the experiment using five different seeds (integers from 0 to 4). The results, which incorporate performance variance, are presented in Table 9 and demonstrate that our approach consistently outperforms the baselines.
Lang. Probability Entropy BoostedProb
en-de 0.142 ± 0.002 0.449 ± 0.030 0.481 ± 0.039
zh-en 0.184 ± 0.004 0.226 ± 0.010 0.312 ± 0.010
Table 9: Performance of QE methods in Pearson correlation across runs with different random seed on NLLB translation.
Appendix E Correlation with ROUGE-L and BertScore
@@
Question Answering Bloomz 560M SciEx en,de -0.002 0.005 0.006
(University Exam) Llama3.2 3B SciEx en,de 0.002 0.228 0.310
Llama3.3 70B SciEx en,de 0.103 0.180 0.180
@@
Question Answering Bloomz 560M SciEx en,de -0.273 0.473 0.497
(University Exam) Llama3.2 3B SciEx en,de 0.102 0.138 0.141
Llama3.3 70B SciEx en,de 0.182 0.204 0.223
@@
Question Answering Bloomz 560M SciEx en,de 0.420 0.411 0.434
(University Exam) Llama3.2 3B SciEx en,de 0.009 0.007 0.015
Llama3.3 70B SciEx en,de 0.005 0.164 0.380
LLM-as-a-Judge Summarization Bloomz 560M XSum en 0.014 0.265 0.287
Llama3.2 3B XSum en -0.039 0.000 0.156
Llama3.3 70B XSum en 0.018 0.130 0.184
Question Answering Bloomz 560M GSM8K en 0.002 0.023 -0.013
(Math) Llama3.2 3B GSM8K en 0.015 -0.204 0.128
Llama3.3 70B GSM8K en 0.024 -0.035 0.053
Question Answering Bloomz 560M SciEx en,de 0.539 0.646 0.654
(University Exam) Llama3.2 3B SciEx en,de -0.030 -0.123 -0.008
Llama3.3 70B SciEx en,de 0.076 0.229 0.305
Table 10: Correlation of different quality scores to different ground truth metrics: Bart Score, RougeL, BertScore, and LLM-as-a-Judge with Qwen2.5 72B.
@@
Appendix F Negative Effect on Very Weak Models
As mentioned in Section 6.4, BoostedProb improves QE performance of output probability for stronger models, but worsens it for weak models. This is somewhat expected, since the motivation of BoostedProb is to improve cases when the model is underconfident. It does not consider the cases when a low-quality model is overconfident and constantly assigns high probability values to the wrong token. To test whether this is truly the cause, we manually look at some output by the worst-performing model, Whisper Tiny, on Chinese-to-English test data. One example is as follows:
Source: "有了它,我们才有了火车、汽车和许多其他交通工具"